quesn 1)the number of tangents tho the curve X^(3/2)+Y^(3/2)=A^(3/2), where the tangents are equally inclined to the axes,is
option a)2 b)1 c)0 d)none of these.

plz send answer to my e-mail id vikash.chandola@yahoo.com

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

x3/2+y3/2=A3/2, differentiating both sides w.r.t. x, (3/2)√x +(3/2)√y(dy/dx)=0

dy/dx=-√(x/y), therefore if tangent is equally inclined to axes, mod(dy/dx)=1, in addition,x3/2+y3/2=A3/2, implies that x,y>0, thus, x=y=A(4)-1/3 is the one and only solution